MICROWAVE AND AB INITIO INVESTIGATIONS OF CHCl_{2}F–OCS AND RELATED HYDROCHLO-ROFLUOROCARBON COMPLEXES

\underline{REBECCA A. PEEBLES}, AMANDA L. STEBER, Eastern Illinois University, Department of Chemistry, 600 Lincoln Ave., Charleston, IL 61920.

The microwave spectrum of the weakly bound dimer of Freon-21 (CHCl_{2}F) with OCS has been investigated by Fouriertransform microwave spectroscopy and ab initio calculations. The two equivalent chlorine atoms in CHCl_{2}F lead to extensive hyperfine splittings in the spectra, and since they do not lie in any of the planes defined by the principal axes of the complex, all three off-diagonal elements of the quadrupole coupling tensor are expected to be nonzero. Initial analysis of the extensive hyperfine splitting from the two chlorine atoms leads to fitted rotational constants that are in excellent agreement with MP2/6-311++G(2d,2p) predictions for a structure with a 2.49 Å C-H...O interaction between the CHCl_{2}F and OCS and a secondary interaction between the chlorine atoms from the CHCl_{2}F and the carbon atom of the OCS (C...C = 3.60 Å; C...Cl = 3.57 Å). Ab initio structures for the CHCl_{2}F–CO_{2}, CHClF_{2}–OCS and CHClF_{2}–CO_{2} complexes will also be examined and comparisons will be made to the analogous series of CH_{n}F_{4−n} (n = 1 − 3) complexes with OCS and CO_{2}.